Find a unit vector u in the direction of v. Verify that llu l = 1. v = (-7, -4)
Answer #1

soine V= -71 - 43 tenit rector af r= -(tê+4;) 555 0-(74+49) JJ6S 11011 - ) 4.5 11811 = 1